Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 81     Entry time: 15 Mar 2023, 01:50
Author: Johannes Moeller 
Group:  
Subject: shift summary 
  • HXRSS delivery, 600 - 700 uJ, 1 bunch, 8.991 keV. Pedestal visibly in HIREX, not checked with MID spectrometer yet.
  • beamline aligned and beam brought to popin imager
  • struggle with jet pumping, to be continued
  • beam brought into experimental hutch, first rough alignment of reference samples and cameras
  • Epix flight path leaked and window re-glued
ELOG V3.1.4-7c3fd00